Product X is a consumer product with a retail price of $9.95. Retailers margins on the product are 40% and wholesalers margins are 8% (based on the selling price). The size of the market is $300,000,000 annually (based on retail sales); product X share (in dollars) of this market is 17.3% The fixed costs involved in manufacturing Product X are $1,400,000 and the variable costs are $0.86 per unit. The advertising budget for Product X is $2,000,000. Miscellane- ous variable costs (e.g., shipping and handling) are $0.04 per unit. Salespeople are paid entirely by a 12% commission based on the manufacturers selling price. Product man agers salary and expenses are $90,000 Assuming that you are the manufacturer, calculate the following: 1. What is the unit margin (contribution) for Product X (in $)? 2. What is Product Xs break-even volume? 3. What market share (based on retail sales) did Product X need to break even? 4. What is Product Xs (annual) net profit? 5. Calculate the increase in sales over the current volume needed to maintain the current profit level if the manufacturer doubles its advertising expenditures 6. Calculate the increase in sales over the current volume needed to maintain the current profit level if the manufacturer lowers its price by 25% Calculate the increase in sales over the current volume needed to maintain the cur- rent profit level if the manufacturer increases the sales forces commission to 15% 7. Calculations Guidelines: When solving problems Round manufacturer, wholesale, and retail prices to the second decimal point (e.g., $9.49) Round sales revenues and net income (profits) to the second decimal point (e.g., $3,625,769.64)
Answer #1
Computaton of Wholesale Price
Wholesaler price = Retail price- retailer margin
= 9.95- 40%*9.95= $5.97
Computaton of Manufacturer Price
Manufacturer price= wholesaler price- wholesaler margin
= 5.97- 8%*5.97= $5.49
Compute Variable ccost per Unit
Variable cost per unit= Manufacturing cost + Miscellaneous cost+ sales commission
=0.86+0.04+12%*5.49=$1.5588
Part-1: Margin contribution per unit= Selling price- Variable cost
= 5.49- 1.5588= $3.9312
Part-2: Break even volume = Fixed costs/ Contribution per unit
= (1400000+2000000+90000)/ 3.9312 = 887789.60 or 887770 Unt
Part-3: Market share required to break even = Break even sales in units*Retail price/ Total market
= 887770*9.95/ 300000000= 0.0294
Part-4: Present sales = 17.3%*300000000=51,900,000
Sales in units= 51,900,000/9.95= 5216080
Net Income = Contribution – Fixed costs
= 5216080*3.9312 - (1400000+2000000+90000)= $17015454